OpenStudy (anonymous):

How do I simplify this and isolate x? 5.2=log(x/25)

OpenStudy (mertsj):

What is the base?

OpenStudy (ksaimouli):

ok so 5.2=log x-log25

OpenStudy (ksaimouli):

@Mertsj the base is 10

OpenStudy (mertsj):

Whose question is this?

OpenStudy (ksaimouli):

we must assume

OpenStudy (mertsj):

Why are you answering instead of the asker?

OpenStudy (ksaimouli):

oops i thought u have the question regarding basses

OpenStudy (kropot72):

\[5.2=\frac{x}{25}\] \[10^{5.2}=\frac{x}{25}\] \[x=10^{5.2}\times 25=you\ can\ calculate\]

OpenStudy (anonymous):

Thank you everyone. kropot72's answer helped me the most.

OpenStudy (kropot72):

You're welcome :)

OpenStudy (kropot72):

Note that in my posting of the method the first equation should be: \[5.2=\log_{} \frac{x}{25}\]